OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 -3x + 2y = 1 5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

add the equations to each other 3x+(-3x) + y+2y = 14+1 0x + 3y 15 3y=15 y=5

OpenStudy (shamil98):

Add the equations together. 3x + y = 14 -3x +2y = 1 ------------ 0 + 3y = 15 divide by 3 y = 5

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

Anytime!